what is the sum of the series 7+22+37...+292?
Found 2 solutions by Boreal, solver91311:
Answer by Boreal(15235)

It is (n/2)(a1+an)
n-1=285/15, the common difference=19
n=20
10*(7+292)=2990

Answer by solver91311(24713)



The sum of an arithmetic series is given by the sum of the first and last terms multiplied by the number of terms, quantity divided by two.

The first term is 7 and the last term is 292. The common difference is 22 minus 7 and 37 minus 22, namely 15. If you subtract 7 from 292 you get 285. 285 divided by 15 is 19, so 15 must have been added to 7 nineteen times to get from 7 to 292. Thus there are 20 terms, the 7 in the beginning and the nineteen 15s.

So:



You can do the arithmetic for yourself

John

My calculator said it, I believe it, that settles it